# JETSCAPE Collaboration

### v1

@@ Abstract
We investigate medium-induced modifications to jet substructure observables that characterize hard splitting patterns in central Pb-Pb collisions at the top energy of the Large Hadron Collider (LHC). Using a multistage Monte Carlo simulation of in-medium jet shower evolution, we explore flavor-dependent medium effects through simulations of inclusive and $\gamma$-tagged jets. The results show that quark jets undergo a non-monotonic modification compared to gluon jets in observables such as the Pb-Pb to $p$-$p$ ratio of the Soft Drop prong angle $r_{g}$, the relative prong transverse momentum $k_{T,g}$ and the groomed mass $m_{g}$ distributions. Due to this non-monotonic modification, $\gamma$-tagged jets, enriched in quark jets, provide surprisingly clear signals of medium-induced structural modifications, distinct from effects dominated by selection bias. This work highlights the potential of hard substructures in $\gamma$-tagged jets as powerful tools for probing the jet-medium interactions in high-energy heavy-ion collisions. All simulations for $\gamma$-tagged jet analyses carried out in this paper used triggered events containing at least one hard photon, which highlights the utility of these observables for future Bayesian analysis.
@@ I Introduction
High-energy heavy-ion collision experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) have established jets as essential tools to probe the quark-gluon plasma (QGP), a strongly coupled, deconfined state of matter created in these collisions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Large transverse momentum partons generated in the initial hard scatterings traverse the QGP medium, undergoing significant interactions that lead to observable modifications compared to proton-proton ($p$-$p$) collisions, where a sufficiently large medium is not formed [21]. Early measurements at RHIC, such as the suppression of high transverse momentum hadrons [22, 23, 24, 25, 26] and the disappearance of di-hadron correlations [27, 28, 29], provided the first experimental evidence of parton energy loss due to the strong interactions with the QGP medium. Recent advances at RHIC and LHC have allowed more detailed studies of reconstructed jets [30, 31, 32, 33], providing deeper insights into the fundamental mechanisms underlying jet-medium interactions, such as medium-induced radiation, scattering with medium constituents, in-medium thermalization, the hydrodynamic evolution of energy deposited by jets, etc.
@@
Recent years have seen a proliferation of observables and analysis techniques referred to as jet substructure, with examples including the jet fragmentation function, jet mass, and jet shape. These reveal medium effects on the internal structure of jets [31, 32, 34, 30, 35, 36, 37, 38, 39, 40, 33, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Advanced grooming techniques [58, 59, 60, 61, 62, 63, 64] extract imprints of hard partonic branchings—occurring early in jet evolution—from final-state hadrons forming reconstructed jets, even though such branchings cannot be directly observed experimentally. Understanding these hard branchings, which form the skeleton of a jet’s internal structure, is crucial to uncovering how jet structures are shaped and modified by medium effects, ultimately leading to $p^{\mathrm{jet}}_{T}$ loss. In vacuum, hard branchings are relatively well-described using perturbative methods, offering robust baselines for studying medium-induced modifications in heavy-ion collisions. Hard branchings have also attracted significant theoretical interest, as they provide a promising avenue to explore jet-medium interactions [65, 66, 67, 68, 69, 70, 71, 64, 72, 73, 74, 75].
@@
Soft Drop [62], one of the most widely used grooming methods in heavy-ion physics, identifies two hard prongs with the largest angular separation within the jet while discarding low-momentum prongs, defining the split as the hardest branching. In our previous study [74], we analyzed the Soft Drop observables $z_{g}$ (momentum fraction) and $r_{g}$ (radial separation) for inclusive jets triggered by their $p^{\mathrm{jet}}_{T}$, using Monte Carlo simulations of in-medium jet shower evolution with the matter$+$lbt multistage model [76, 77, 78, 79, 80, 81, 82, 20] within the jetscape framework [83, 84, 85, 86, 87, 88, 89, 90, 74, 91, 21, 92, 93]. The $z_{g}$ distribution showed negligible medium-induced modifications, consistent with experiments, while the $r_{g}$ distribution exhibited suppression that increases monotonically with $r_{g}$. This monotonic behavior was reproduced by incorporating modified coherence effects [94, 89, 95], where high-virtuality partons, due to finer resolution, perceive the medium as dilute and experience reduced interactions. As a result, jets with larger $r_{g}$ were strongly suppressed, with no trace of medium-induced broadening in the hard partonic branchings.
@@
These findings suggest that the suppression of the $r_{g}$ distribution may arise primarily from selection bias in inclusive jet analyses, rather than direct modifications to hard branching structures: Jets with large-angle hard branchings are more likely to lose soft constituents outside the jet cone, leading to a preference for triggering on jets with smaller angular separations and reduced energy loss. Thus, inclusive jet measurements are less sensitive to medium-induced changes in hard branching structures, underscoring the need for targeted observables to probe these effects.
@@
To investigate medium-induced modifications to hard branchings in detail, mitigating the selection bias inherent in inclusive jet analyses is essential. This can be achieved using $\gamma$-tagged or $Z$-tagged jets, produced in photon-jet or $Z$-boson-jet pair production events, often referred to as golden channels. Photons and $Z$ bosons do not interact strongly with the QGP medium. Thus, photons and $Z$ bosons that are pair-produced with a jet in the initial hard scattering are measured with transverse momenta ($p_{T}$) that closely approximate the initial $p_{T}$ of the partons generating the jets, at leading order. Triggering on the $p_{T}$ of the photon or $Z$ boson, instead of the jet, selects jets with similar initial $p_{T}$, independent of $p^{\mathrm{jet}}_{T}$ loss, reducing selection bias and enabling cleaner investigations of medium effects on jet structures. Additionally, partons paired with photons or $Z$ bosons are predominantly quarks, minimizing flavor-dependent selection bias. Comparing these tagged jets with inclusive jets allows systematic studies of flavor-dependent effects.
@@
In this paper, we investigate $\gamma$-tagged jets as a means to mitigate the selection bias effect, enabling a clearer examination of medium modifications to a subset of Soft Drop observables that characterize the hard substructure of jets. Building on our previous work [89, 90, 74], which established a wealth of benchmark results, we employ the matter$+$lbt multistage model within the jetscape framework to provide predictions under realistic high-energy heavy-ion collision configurations. We demonstrate that fixing the photon $p_{T}$ and triggering on associated jets, even those with lower $p^{\mathrm{jet}}_{T}$, eliminates the suppression of jets with broad hard splittings observed in inclusive jet measurements. This confirms that the suppression in inclusive jets arises from selection bias rather than intrinsic modifications to the hard-splitting structure. Moreover, we show that $\gamma$-tagged jets exhibit pronounced broadening of splittings compared to inclusive jets, driven by their quark-jet dominance. With simplified simulations, we reveal that quark jets are more susceptible to medium-induced modifications, particularly during low-virtuality evolution, leading to significant modifications in their hard branching structure. These results establish $\gamma$-tagged jets, and similarly $Z$-tagged jets, as powerful tools for revealing clear signals of medium modifications to the hard branchings in jets through their substructure observables.
@@
The paper is organized as follows. Section II outlines the simulation framework and methodologies employed, including the multistage matter$+$lbt model within the jetscape framework. In Sec. III, we detail the analysis procedures for $\gamma$-tagged jets as well as inclusive jets, including Soft Drop grooming. Section IV presents the results from the jetscape simulations, highlighting the flavor-dependent medium effects and selection biases on $r_{g}$, $k_{T,g}$, and $m_{g}$ distributions, as well as the role of $\gamma$-tagged jets in revealing intrinsic medium modifications. Subsequently, in Sec. V, we compare the results from the same event sets with available experimental data from Pb-Pb collisions at the LHC. Finally, Sec. VI concludes with a summary of our findings.
@@ II Model
In this paper, the jetscape code package is utilized, which offers a flexible framework for modular integration in Monte Carlo event generation for heavy-ion collisions. We specifically employ our current default configuration with the matter$+$lbt setup referred to as the JETSCAPEv3.5 AA22 tune, as detailed in our previous work [89, 90, 74], to perform simulations of jet events. All parameters in the tune are chosen to fit only the $R_{AA}$s for single high-$p_{T}$ particles and reconstructed jets in Ref. [89] and are not retuned for any other observables, including all the observables presented in this paper. In this section, we provide a brief overview of the components that constitute our simulation setup. For readers interested in delving deeper into the specific physical elements integrated into the multistage matter$+$lbt simulation within the jetscape framework, we direct them to Refs. [89, 90, 74]. Additional insights into the software components of the jetscape framework are detailed in Ref. [83]. The foundational concept concerning the multistage description of jet evolution within the medium is explained in Ref. [82].
@@ II.1 Overview
For the efficient generation of jet events, the jetscape framework provides an option to embed a single hard scattering event from a nucleon-nucleon collision into a heavy-ion collision event using a pre-generated space-time background profile of the QGP medium.
In the JETSCAPEv3.5 AA22 tune, we employ the background medium profile obtained from event-by-event calculations of $(2+1)$-dimensional [$(2+1)$-D] free-streaming pre-equilibrium evolution [96], followed by viscous hydrodynamic evolution using $(2+1)$-D vishnu [97] and hadronic scattering and decay simulated by urqmd [98, 99], with fluctuating initial conditions from trento [100]. Here the best-fit parametrization, determined through maximum a posteriori (MAP) from Bayesian methods [101] applied to observables measured at the LHC, is used. In jet simulations, high-energy partons are generated according to the hard process settings described later in Subsection II.2, and then these partons evolve into parton showers. In the multistage setup of matter$+$lbt, these partons are initially subjected to virtuality-ordered splitting while incorporating medium effects in the matter module [77, 76]. The partons then undergo successive parton splittings, rapidly reducing their virtuality. The description of the splittings for a parton in the matter module is terminated when its virtuality becomes sufficiently small, approximately on par with the accumulated transverse momentum gain via scatterings in the medium. This termination is attributed to the applicable limit of a model that relies on virtuality as the primary mechanism for branching. The partons with reduced virtuality are then transferred to the lbt module [102, 78, 79], which simulates elastic and inelastic scatterings with medium constituents based on kinetic theory with the on-shell approximation.
@@ II.2 Initial hard process
In this paper, we compare our results from the jetscape simulations with several different settings for the initial hard process generation. Below, we explain each setting. With any setting, the geometric position in the transverse plane at mid-space-time rapidity ($\eta_{s}=(1/2)\ln[(t+z)/(t-z)]=0$), where the hard process occurs is determined by sampling the $N_{\mathrm{coll}}$ distribution calculated using the trento initial condition module for the Pb-Pb collisions.
@@ II.2.2 PythiaGun
For the generation of realistic jet events in hadron-hadron collisions, the jetscape framework provides the pythiagun module, which serves as a wrapper for pythia[132]. Within the pythiagun module, utilizing the functionalities of pythia 8, hard scatterings are generated based on leading-order perturbation calculations. Subsequently, final state radiation (FSR) is disabled by default, and the resulting hard partons are directly passed to a module responsible for the evolution of highly virtual partons, matter in this study, to facilitate the development of jet showers. Simultaneously, when initial state radiation (ISR) or multiparton interaction (MPI) is turned on, the partons produced through such a process are also passed to the jet shower evolution. Throughout this study, both ISR and MPI are always on for initial scattering generation using pythiagun.
@@
In the actual experimental analysis, the isolation cut mentioned later in Subsec. III.2 is applied to select $\gamma$-tagged jets to enhance the leading-order contributions in initial hard scatterings involving prompt photon production. However, in practice, perfect extraction is not achievable, and contributions beyond leading-order prompt photons unavoidably seep in. For example, if a photon is emitted at a large angle from a parton jet shower, it may pass the isolation cut requirement.

### v2

@@ Abstract
We investigate medium-induced modifications to jet substructure observables that characterize hard components in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. Using a multistage Monte Carlo simulation of in-medium jet shower evolution, we explore flavor-dependent medium effects through simulations of inclusive and $\gamma$-tagged jets. The results show that quark jets undergo a nonmonotonic modification compared with gluon jets in observables such as the Pb-Pb to $p$-$p$ ratio of the soft drop prong angle $r_{g}$, the relative prong transverse momentum $k_{T,g}$, and the groomed mass $m_{g}$ distributions. Due to this nonmonotonic modification, $\gamma$-tagged jets, enriched in quark jets, provide surprisingly clear signals of medium-induced structural modifications, distinct from effects dominated by selection bias. Further systematic studies demonstrate that these effects are dominated by recoil medium response. This work highlights the potential of hard substructures in $\gamma$-tagged jets as powerful tools for probing the jet-medium interactions in high-energy heavy-ion collisions, in particular by enabling detailed investigations of jet-medium parton scatterings via their associated medium response. All simulations for $\gamma$-tagged jet analyses carried out in this paper used triggered events containing at least one hard photon, which highlights the utility of these observables for future Bayesian analysis.
@@ I Introduction
High-energy heavy-ion collision experiments at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC) have established jets as essential tools to probe the quark-gluon plasma (QGP), a strongly coupled, deconfined state of matter created in these collisions [38, 30, 34, 33, 145, 80, 78, 79, 139, 140, 77, 137, 99, 31, 32, 98, 39, 114, 50, 48]. Large transverse momentum partons generated in the initial hard scatterings traverse the QGP medium, undergoing significant interactions that lead to observable modifications compared with proton-proton ($p$-$p$) collisions, where a sufficiently large medium is not formed [128]. Early measurements at RHIC, such as the suppression of high transverse momentum hadrons [24, 28, 27, 25, 21] and the disappearance of di-hadron correlations [26, 22, 23], provided the first experimental evidence of parton energy loss due to the strong interactions with the QGP medium. Recent advances at RHIC and LHC have allowed more detailed studies of reconstructed jets [13, 7, 59, 19], providing deeper insights into the fundamental mechanisms underlying jet-medium interactions, such as medium-induced radiation, scattering with medium constituents, in-medium thermalization, the hydrodynamic evolution of energy deposited by jets, etc.
@@
Recent years have seen a proliferation of observables and analysis techniques referred to as jet substructure, with examples including the jet fragmentation function, jet mass, and jet shape. These reveal medium effects on the internal structure of jets [7, 59, 8, 13, 10, 85, 20, 6, 15, 125, 19, 61, 84, 122, 124, 123, 14, 126, 60, 9, 3, 4, 5, 11, 29, 81, 18, 17]. Advanced grooming techniques [41, 86, 71, 67, 91, 92, 102] extract imprints of hard partonic branchings from final-state hadrons forming reconstructed jets, even though such branchings cannot be directly observed experimentally. Understanding these hard branchings, which form the skeleton of a jet’s internal structure, is crucial to uncovering how jet structures are formed and modified by medium effects, such as modifications of the radiation pattern or scatterings with medium constituents, ultimately leading to $p^{\mathrm{jet}}_{T}$ loss. In vacuum, hard branchings are relatively well-described using perturbative methods, offering robust baselines for studying medium-induced modifications in heavy-ion collisions. Hard branchings have also attracted significant theoretical interest, as they provide a promising avenue to explore jet-medium interactions [64, 103, 57, 104, 51, 116, 54, 102, 55, 56, 131, 65].
Here, it is important to note that, especially in the presence of medium effects, the notion of a hard branching extracted by grooming analyses does not necessarily correspond to an explicit parton radiation. The reconstructed hard branching may instead be strongly affected by recoil particles in scatterings with medium constituents [104], contributions from hadronization processes, or other effects that, in principle, cannot be disentangled experimentally. Accordingly, throughout this paper, the term “hard branching” denotes the splitting identified by the grooming analysis, whereas the hard substructure of the jet refers to observables characterizing the distribution of the reconstructed prongs associated with that branching.
@@
Soft drop [91], one of the most widely used grooming methods in heavy-ion physics, identifies the two hard prongs with the largest angular separation within a jet while discarding low-momentum prongs, thereby defining the hard branching. In our previous study [131], we analyzed the soft-drop observables $z_{g}$ (momentum fraction) and $r_{g}$ (radial separation) for inclusive jets triggered by their $p^{\mathrm{jet}}_{T}$, using Monte Carlo simulations of in-medium jet shower evolution with the matter++lbt multistage model [49, 100, 83, 47, 82, 94, 45, 48] within the jetscape framework [112, 87, 73, 72, 46, 74, 88, 75, 131, 76, 128, 69, 121]. The $z_{g}$ distribution showed negligible medium-induced modifications, consistent with experiments, while the $r_{g}$ distribution exhibited suppression that increases monotonically with $r_{g}$. This monotonic behavior was reproduced by incorporating modified coherence effects [89, 88, 90], where high-virtuality partons, due to finer resolution, perceive the medium as dilute and experience reduced interactions. As a result, jets with larger $r_{g}$ were strongly suppressed, with no trace of medium effects on the actual structure of hard components associated with the parton shower.
@@
These findings suggest that the suppression of the $r_{g}$ distribution may arise primarily from selection bias in inclusive jet analyses, rather than from direct structural modifications of soft-drop-reconstructed hard branchings. Jets with large-angle hard branchings are more likely to have lost soft constituents outside the jet cone, leading to a bias toward triggering on jets with smaller angular separations and reduced overall energy. Thus, inclusive jet measurements are less sensitive to medium effects, causing the structural modification of the parton shower, underscoring the need for targeted observables to probe these effects.
@@
To investigate medium effects in detail through soft-drop-reconstructed jet substructure, it is essential to mitigate the selection bias inherent in inclusive jet analyses. This can be achieved using $\gamma$-tagged or $Z$-tagged jets, produced in photon-jet or $Z$-boson-jet pair production events, often referred to as “golden channels.” Photons and $Z$ bosons do not interact strongly with the QGP medium. Thus, photons and $Z$ bosons that are pair-produced with a jet in the initial hard scattering are measured with transverse momenta ($p_{T}$) that closely approximate the initial $p_{T}$ of the partons generating the jets, at leading order. Triggering on the $p_{T}$ of the photon or $Z$ boson, instead of the jet, selects jets with similar initial $p_{T}$, independent of $p^{\mathrm{jet}}_{T}$ loss, reducing selection bias and enabling cleaner investigations of medium effects on jet structures. Additionally, partons paired with photons or $Z$ bosons are predominantly quarks, minimizing flavor-dependent selection bias. Comparing these tagged jets with inclusive jets allows systematic studies of flavor-dependent effects.
@@
In this paper, we investigate $\gamma$-tagged jets as a means to mitigate the selection bias effect, enabling a clearer examination of medium modifications to a subset of soft-drop observables that characterize the hard substructure of jets. Building on our previous work [88, 75, 131], which established a wealth of benchmark results, we employ the matter++lbt multistage model within the jetscape framework to provide predictions under realistic high-energy heavy-ion collision configurations. We demonstrate that fixing the photon $p_{T}$ and triggering on associated jets, even those with lower $p^{\mathrm{jet}}_{T}$, eliminates the suppression of jets with broad hard branchings observed in inclusive jet measurements. This confirms that the suppression of inclusive jets arises from selection bias rather than from intrinsic structural modifications of the hard component reconstructed as part of a jet. Moreover, we show that $\gamma$-tagged jets exhibit pronounced broadening of soft-drop-reconstructed hard branchings compared with inclusive jets, driven by their quark-jet dominance. With simplified simulations, we reveal that quark jets are more susceptible to medium-induced modifications, particularly during low-virtuality evolution, leading to significant modifications in their soft-drop-reconstructed hard branching structure. These results establish $\gamma$-tagged jets, and similarly $Z$-tagged jets, as powerful tools for revealing clear signals of direct medium effects on the hard component reconstructed as part of a jet.
@@
The paper is organized as follows. Section II outlines the simulation framework and methodologies employed, including the multistage matter++lbt model within the jetscape framework. In Sec. III, we detail the analysis procedures for $\gamma$-tagged jets as well as inclusive jets, including soft-drop grooming. Section IV presents the results from the jetscape simulations, highlighting the flavor-dependent medium effects and selection biases on $r_{g}$, $k_{T,g}$, and $m_{g}$ distributions, as well as the role of $\gamma$-tagged jets in revealing intrinsic medium modifications. Subsequently, in Sec. V, we compare the results from the same event sets with available experimental data from Pb-Pb collisions at the LHC. Finally, Sec. VI concludes with a summary of our findings.
@@ II Model
In this paper, the jetscape code package is utilized, which offers a flexible framework for modular integration in Monte Carlo event generation for heavy-ion collisions. We specifically employ our current default configuration with the matter++lbt setup referred to as the JETSCAPEv3.5 AA22 tune, as detailed in our previous work [88, 75, 131], to perform simulations of jet events. All parameters in the tune are chosen to fit only the $R_{AA}$ s for single high-$p_{T}$ particles and reconstructed jets in Ref. [88] and are not retuned for any other observables, including all the observables presented in this paper. In this section, we provide a brief overview of the components that constitute our simulation setup. For readers interested in delving deeper into the specific physical elements integrated into the multistage matter++lbt simulation within the jetscape framework, we direct them to Refs. [88, 75, 131]. Additional insights into the software components of the jetscape framework are detailed in Ref. [112]. The foundational concept concerning the multistage description of jet evolution within the medium is explained in Ref. [45].
@@ II.1 Overview
For the efficient generation of jet events, the jetscape framework provides an option to embed a single hard scattering event from a nucleon-nucleon collision into a heavy-ion collision event using a pre-generated space-time background profile of the QGP medium. In the JETSCAPEv3.5 AA22 tune, we employ the background medium profile obtained from event-by-event calculations of $(2+1)$-dimensional [$(2+1)$-D] free-streaming pre-equilibrium evolution [93], followed by viscous hydrodynamic evolution using $(2+1)$-D vishnu [120] and hadronic scattering and decay simulated by urqmd [35, 40], with fluctuating initial conditions from trento [105]. Here, the best-fit parametrization, determined through maximum a posteriori (MAP) from Bayesian methods [36] applied to observables measured at the LHC, is used.
In jet simulations, high-energy partons are generated according to the hard process settings described later in Sec. II.2, and then these partons evolve into parton showers. In the multistage setup of matter++lbt, partons are initially subjected to splittings driven primarily by their large virtualities, with minor medium effects incorporated in the matter module [100, 49]. In matter, the splittings are simulated in a virtuality-ordered manner, while the space-time trajectory of the shower is constructed by estimating the formation time via the uncertainty principle and sequentially determining the splitting locations. The description of the splittings for a parton in the matter module is terminated when its virtuality becomes sufficiently small, approximately on par with the accumulated transverse momentum gain via scatterings in the medium. This termination is attributed to the applicable limit of a model that relies on virtuality as the primary mechanism for partonic splittings. The partons with reduced virtuality are then transferred to the lbt module [138, 83, 47], which simulates elastic and inelastic scatterings with medium constituents based on kinetic theory with the on-shell approximation.
@@
In this paper, to investigate the impact of medium response effects via recoils and holes, we also present results from calculations in which both recoils and holes are excluded. For this setup, the parton shower evolution is carried out identically to the default simulations with recoils and holes included; however, at the hadronization stage, all recoil and hole partons are discarded and do not contribute to final-state particles. Since all $s$-, $t$-, and $u$-channels contribute to the jet-medium scatterings, it is in principle impossible to distinguish whether an outgoing parton originates from the jet shower or from the medium. In practice, the scattered parton with the smaller energy is tagged as a recoil. It should also be noted that calculations without recoils and holes correspond to an unphysical setup that explicitly violates energy-momentum conservation and are intended solely for theoretical assessments.
@@ II.2 Initial hard process
In this paper, we compare our results from the jetscape simulations with several different settings for the initial hard process generation. Below, we explain each setting. With any setting, the geometric position in the transverse plane at mid-space-time rapidity $\eta_{s}=(1/2)\ln[(t+z)/(t-z)]=0$], where the hard process occurs, is determined by sampling the $N_{\mathrm{coll}}$ distribution calculated using the trento initial condition module for the Pb-Pb collisions.
@@ II.2.2 PythiaGun
For the generation of realistic jet events in hadron-hadron collisions, the jetscape framework provides the pythiagun module, which serves as a wrapper for pythia[127]. Within the pythiagun module, utilizing the functionalities of pythia 8, hard scatterings are generated based on leading-order perturbation calculations. Subsequently, final-state radiation (FSR) is disabled by default, and the resulting hard partons are directly passed to a module responsible for the evolution of highly virtual partons, matter in this study, to facilitate the development of jet showers. Simultaneously, when initial-state radiation (ISR) or multiparton interaction (MPI) is turned on, the partons produced through such a process are also passed to the jet shower evolution. Throughout this study, both ISR and MPI are always on for initial scattering generation using pythiagun.
@@
In the actual experimental analysis, the isolation cut mentioned later in Sec. III.2 is applied to select $\gamma$-tagged jets to enhance the leading-order contributions in initial hard scatterings involving prompt photon production. However, in practice, perfect extraction is not achievable, and contributions beyond leading-order prompt photons unavoidably seep in. For example, if a photon is emitted at a large angle from a parton jet shower, it may pass the isolation cut requirement.